OpenStudy (anonymous):

What volume of 1.25 M sodium hydroxide (NaOH) solution can be prepared using 60.0 g of sodium hydroxide? (The mass of one mole of NaOH is 40.0 g.) Answer 0.667 L 0.833 L 1.20 L 1.88 L

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

I'm sorry, I have to leave soon, but maybe @ganeshie8 could help you or know someone who is online and can. All the best.

OpenStudy (abmon98):

Number of moles=Mass(g)/Molar Mass(g/mol) Number of moles=Concentration(mol/dm^3)*Volume(dm^3) 60/40=3/2 moles of NaOH (1.5)/1.25

OpenStudy (anonymous):

1.2 is correct then? Also, thnx for your response @abb0t :)

OpenStudy (anonymous):

I had already selected 1.2, but wasn't sure..

OpenStudy (abmon98):

yes its 1.2

OpenStudy (anonymous):

gratzi!! :D